# Getting Bored of Cyberwar: Exploring the Role of Low-level Cybercrime Actors in the Russia-Ukraine Conflict

### v6

@@ Acknowledgments
We are grateful to the anonymous reviewers and our colleagues at the Cambridge Cybercrime Centre for their insightful and constructive feedback. This work was supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 949127). The conclusions and findings demonstrated in this study are those of the authors and do not necessarily reflect the opinions of the funders.
@@ Appendix A Ethical Considerations
This work is presented objectively to minimise risk to researchers. The collection, sharing, and analysis of web defacement, amplified DDoS attacks, and Telegram chats have been formally approved by the Department of Computer Science & Technology’s Research Ethics Committee. We do not attempt to gather private data; only publicly accessible data are collected. A 2022 U.S. court ruling suggests scraping public data is legal (TechCrunch, 2022); our scraper does not overload websites. The amplified DDoS attack honeypots absorb attack packets without relaying them, thus reducing harm to victims.
Studying an ongoing conflict may harm individuals whose attacks are reported, while researchers might face retaliation from attackers due to leaking insights into their activities and community. To avoid potential harms, our experiments operate ethically and collectively, only presenting aggregated findings without identifying individuals. We did not ask for consent from Telegram users or web defacers when using scraped data, as sending thousands of messages would be impractical. We believe they are aware that their content posted online are publicly visible. This approach accords with the British Society of Criminology’s Statement on Ethics (British Society of Criminology, 2015).
@@ Appendix B Data Licensing
We have robust procedures and long experience in licensing our data in various jurisdictions. Our quantitative data and analysis code are available for academic researchers under a license agreement with the Cambridge Cybercrime Centre to prevent misuse and to ensure the data will be treated ethically, as access to sensitive data might risk both researchers and the actors involved (Warford et al., 2022).
@@ Appendix C Determining Attack Geolocation
For example, a ‘.ru’ website is supposed to be Russian, but it might be physically hosted in Vietnam, operated by a person living in Hong Kong, while proxied through Cloudflare with an IP address in the US. Relying on only one aspect might be risky, as both IP and domain can lie. We use data fusion to enhance the accuracy, prioritising: (1) top-level domain; (2) IP geolocation at collection time (MaxMind GeoIP2222 GeoIP2 is freely accessible at https://maxmind.com/. It offers both free and paid licenses, with the paid one being slightly more accurate and up-to-date. It claims to provide over 99.8% country-level and over 60% city-level accuracy, yet that varies from country to country e.g., 79% for Russia and 65% for Ukraine, within a 250km radius. for web defacement, and a database we maintain based on Regional Internet Registry data for DDoS attacks; (3) geolocation of the AS hosting the IP address. If a website’s IP address belongs to a CDN, its geolocation is determined solely by ccTLD, as any geolocation of IP address or ASN will be unreliable.
The top three CDNs are Cloudflare, Amazon Web Services, and Akamai, serving around 89% of customers (Intricately Market Analysts, 2020). We ignore the trivial market shares of their competitors, but count DDoS-Guard as it is Russian-based, which may affect the measurement. We expect the four cover nearly 90% of customers. We found 4.87% of defacements are hosed on these CDNs by 14 262 prefixes as of the writing date: 1 698 of Cloudflare; 7 483 of Amazon Webservice; 5 056 of Akamai; and 25 of DDoS-Guard (these prefixes and AS number mappings are collected on Hurricane Electric Internet Services). For defacements, we prefer ccTLDs over IP geolocation as attackers often target sites by massively scanning domain ccTLDs, such as ‘.ru’ and ‘.ua’, rather than checking whether IP addresses are hosted in those countries.
@@ Appendix D Defacement Submission Process
The defacement submission is mostly automatic: users specify a ‘notifier’, team, defaced URL, vulnerability types, and hacking incentives. At that point, a record is made with details of the compromised system, its IP address and location, and a snapshot of the defaced page (often consisting of messages that may include political and ideological propaganda (Banerjee et al., 2022)). Messages can be hidden using identical font colours as the background, but are detectable through HTMLs. New reports are kept away from the dashboard until being verified by staff or bots. Although ‘notifier’ can be arbitrarily entered, defacers are incentivised to use consistent handles to cultivate fame and reputation; we thus consider ‘notifier’ to be reliable enough to differentiate them. Snapshots of defaced sites, including messages left, are highly reliable as they are captured at reporting time.
Appendix E Web Defacement Collection
Data completeness and reliability are critical for longitudinally measurement. Scraping complete defacement archives at scale, especially Zone-H, is non-trivial, and was not guaranteed in prior work. Some purchased the data (Maggi et al., 2018), but it is not sustainable and is ethically questionable. This is challenging as (1) Zone-H uses Captcha to prevent bots, (2) its dashboard sets a limit of 50 pages where older data is hidden, and (3) on-hold records may not appear promptly, leading to potential misses. The only way to get a complete scrape is iterating all submission IDs, with IDs of valid and invalid reports often mixed. This issue, along with IP blacklisting and bot prevention, generates non-trivial workload. We responded by (1) developing an efficient text Captcha solver for Zone-H using image-processing techniques, (2) routing our scraper through multiple proxies, and (3) iterating all submission IDs in turn. Raw data is stored in a database to avoid unnecessary future requests.
@@ Appendix F Validating On-hold Defacements
How defacement submissions are validated is not clearly stated. While Zone-H reports are kept on hold until being manually verified by staff, Zone-Xsec, Defacer-Pro, and Haxor-ID use automatic validation, insisting messages left on the defaced pages contain keywords linked to hacking activities (e.g., ‘Hacked by Me’). Defacers may game the system by putting comments on blogs, or submitting search queries (e.g., ?search=‘Hacked by Me’), which occasionally get through automatic sanitisation but our further validation excludes them. Manual staff review on Zone-H may be slow, while automatic verification of the others is error-prone. Unverified records may be kept on-hold forever, leading to incomplete data. Consequently, collecting only defacements shown in the dashboard is inadequate, making a complete dataset challenging to gather. To enhance data completeness and reliability, a semi-automatic validation is performed to check if on-hold reports are in fact valid.
@@
Our our strategy is shown in Algorithm 1. First, reports verified by archives are considered valid. Second, messages on the defaced pages of on-hold submissions are used to decide the validity, as defacers often leave signatures for reputation e.g., ‘Hacked by CoolHacker’. If messages include defacers’ handles and specifically contain common hacking terms: ‘hacked by’, ‘h4ck3d by’, ‘h4cked by’, ‘p4wn3d by’, ‘pwn3d by’, ‘pwnd by’, ‘pwned by’, ‘pwndz by’, ‘owned by’, ‘own3d by’, ‘touched by’, and ‘kissed by’, we consider them to be valid e.g., a message ‘This website was hacked, contact me t.me/coolhacker’ posted by a notifier ‘CoolHacker’ is considered valid. This method is looser than an exact comparison with ‘Hacked by CoolHacker’, but is still highly accurate; 100 randomly checked samples were all correct. Third, the remaining submissions are manually validated by looking for defacers’ signature; some are obvious and some are complicated. Candidates are grouped by normalised handles and messages (redundant spaces removed), then for each, 10 most similar validated defacements are suggested to the annotator. Levenshtein distance is used to estimate the similarity between two messages, which is helpful as messages are often slightly modified from existing templates. If no message is found (instead images, iframes, or javascript), or leftover signatures cannot be spotted, a web browser opens the defaced page and assist annotators.
@@
This assistance effectively reduces the annotator’s effort. One challenge is the redirects to a defacer’s page, as this can be modified dynamically; when the defacer’s page is down, the submission points to a non-existent site, but a careful check could reveal evidence of the defacers. We also consider a site is touched if its content is unchanged but the page title is modified to indicate hacking activities. We ignore cases that lack evidence to ensure that those flagged ‘valid’ are indeed valid. We do not use complex machine learning techniques as message texts contain lots of noise; given a small number of samples (around 10k), machine learning is not more effective than a rule-based approach. Sometimes defacements appeared to be already verified at the time of collection, but became invalid afterwards; we re-validate them months after collection to make sure their status has been finalised by the archives.
@@ Appendix G Unifying Defacements and Defacers
We hashed then unified defacements based on reporting dates, original defacer handles, root victim domains, and message left on the defaced page. Including reporting dates may be an issue if defacers resubmit to other archives after a few days; but excluding them may lead to spotting matched submissions on different days due to repeat victimisation. We also unified defacers across all archives, as users tend to pick similar pseudonyms on different platforms (Goga et al., 2013). As the unification needs to be accurate and the number of unique defacers does not exceed a few thousand, machine learning is not appropriate. We instead used a semi-automated approach combining automated handle similarity analysis with manual review. First, similar handle pairs are extracted using Levenshtein distance, which is set to not exceed 25% of the handles’ length. Then ten messages left by defacers in each pair are sampled to assist the annotation; the pair is unified under a single nickname if their messages are semantically closed enough, based on message patterns, stylometry, synonyms, handle appearances, typos, team, nationality, messages’ semantics, language, and the handle rarity (rare ones like ‘cj2ks’ are more likely to be used by a single person, while common ones like ‘glory’ are more likely to be shared by multiple individuals (Liu et al., 2013)). Many handles leave similar messages on different archives, while different typos occur e.g., missing characters, order of characters, case-sensitive. Messages left are diverse: some are identical, some are relatively similar, some are distinct and contain the defacers’ name, and some come with phone numbers. We only confirm when having sufficient evidence, uncertain pairs are left unmatched.

### v7

@@ Acknowledgments
We thank the anonymous reviewers. This work is supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme, grant No 949127.
@@ Appendix A Ethical Considerations
This work is presented objectively to minimise risk to researchers. The collection, sharing, and analysis of web defacement, amplified DDoS attacks, and Telegram chats have been formally approved by the Department of Computer Science & Technology’s Research Ethics Committee. We do not attempt to gather private data; only publicly accessible data are collected. A 2022 US court ruling suggests scraping public data is legal (TechCrunch, 2022); our scraper does not overload websites. The amplified DDoS attack honeypots absorb attack packets without relaying them, thus reducing harm to victims.
Studying an ongoing conflict may harm individuals whose attacks are reported, while researchers might face retaliation from attackers due to leaking insights into their activities and community. To avoid potential harm, our experiments operate ethically and collectively, only presenting aggregated findings without identifying individuals. We did not ask for consent from Telegram users or web defacers when using scraped data, as sending thousands of messages would be impractical. We believe they are aware that content posted online will be publicly visible. This approach accords with the British Society of Criminology’s Statement on Ethics (British Society of Criminology, 2015).
@@ Appendix B Data Licensing
We have robust procedures and long experience in licensing our data in various jurisdictions. Our quantitative data, analysis scripts, and scrapers are available for academic researchers under a license agreement with the Cambridge Cybercrime Centre to prevent misuse and to ensure the data will be treated ethically, as access to sensitive data might risk researchers and the actors involved (Doerfler et al., 2021; Warford et al., 2022).
@@ Appendix C Determining Attack Geolocation
For example, a ‘.ru’ website is supposed to be Russian, but it might be physically hosted in Vietnam, operated by a person living in Hong Kong, while proxied through Cloudflare with an IP address in the US. Relying on only one aspect might be risky, as both IP and domain can lie. We use data fusion to enhance the accuracy, prioritising: (1) top-level domain; (2) IP geolocation at collection time (MaxMind GeoIP2222 GeoIP2 is freely accessible at https://maxmind.com/. It offers both free and paid licenses, with the paid one being slightly more accurate and up-to-date. It claims to provide over 99.8% country-level and over 60% city-level accuracy, yet that varies from country to country e.g., 79% for Russia and 65% for Ukraine, within a 250km radius. for web defacements, and a database we maintain based on Regional Internet Registry data for DDoS attacks; and (3) geolocation of the AS hosting the IP address. If a website’s IP address belongs to a CDN, its geolocation is determined solely by ccTLD, as any geolocation of IP address or ASN will be unreliable.
The top three CDNs are Cloudflare, Amazon Web Services, and Akamai, serving around 89% of customers (Intricately Market Analysts, 2020). We ignore the trivial market shares of their competitors but count DDoS-Guard as it is Russian-based, which may affect the measurement. We expect the four to cover over 90% of customers. We found 4.87% of defacements are hosted on these CDNs by 14 262 prefixes as of the writing date: 1 698 of Cloudflare; 7 483 of Amazon Webservice; 5 056 of Akamai; and 25 of DDoS-Guard (these prefixes and AS number mappings are collected on Hurricane Electric Internet Services). For defacements, we prefer ccTLDs over IP geolocation as attackers often target sites by massively scanning domain ccTLDs, such as ‘.ru’ and ‘.ua’, rather than checking whether IP addresses are hosted in those countries.
@@ Appendix D Defacement Submission Process
The defacement submission is mostly automatic: users specify a ‘notifier’, team, defaced URL, vulnerability types, and hacking incentives. At that point, a record is made with details of the compromised system, its IP address and location, and a snapshot of the defaced page (often consisting of messages that may include political and ideological propaganda (Banerjee et al., 2022)). Messages can be hidden using identical font colours as the background, but are detectable through HTML. New reports are kept away from the dashboard until being verified by staff or bots. Although ‘notifier’ can be arbitrarily entered, defacers are incentivised to use consistent handles to cultivate fame and reputation; we thus consider ‘notifier’ to be reliable enough to differentiate them. The defacement snapshots, including messages left, are highly reliable as they are captured at reporting time.
Appendix E Website Defacement Collection
Data completeness and reliability are critical for longitudinal measurement. Scraping complete defacement archives at scale, especially Zone-H, is non-trivial, and was not guaranteed in prior work. Some purchased the data (Maggi et al., 2018), but it is not sustainable and is ethically questionable. This is challenging as (1) Zone-H uses Captcha to prevent bots, (2) its dashboard sets a limit of 50 pages where older data is hidden, and (3) on-hold records may not appear promptly, leading to potential misses. The only way to get a complete scrape is by iterating all submission IDs, with IDs of valid and invalid reports often mixed. This issue, along with IP blacklisting and bot prevention, generates a non-trivial workload. We responded by (1) developing an efficient text Captcha solver for Zone-H using image-processing techniques, (2) routing our scraper through multiple proxies, and (3) iterating all submission IDs in turn. Raw data is stored in a database to avoid unnecessary future requests.
@@ Appendix F Validating On-hold Defacements
How defacement reports are validated is not clearly stated. Zone-H reports are kept on hold until being manually verified by staff; Zone-Xsec, Defacer-Pro, and Haxor-ID use automatic validation, insisting messages left on defaced pages contain keywords linked to hacking activities (e.g., ‘Hacked by ABC’); while it is unclear for OwnzYou. Defacers may game the system by putting comments on blogs, or submitting search queries (e.g., ?search=‘Hacked by ABC’), which occasionally get through automatic sanitisation but our further validation excludes them. Manual staff review on Zone-H may be slow, while automatic verification of the others is error-prone. Unverified records may be kept on hold forever, leading to incomplete data. Consequently, collecting only defacements shown in the dashboard is inadequate, making a complete dataset challenging to gather. To enhance data completeness and reliability, a semi-automatic validation is performed to check if on-hold reports are in fact valid.
@@
Our strategy is shown in Algorithm 1. First, reports verified by archives are considered valid. Second, messages on the defaced pages of on-hold submissions are used to decide the validity, as defacers often leave signatures for reputation e.g., ‘Hacked by CoolHacker’. If messages include defacer handles and specifically contain common hacking terms: ‘hacked by’, ‘h4ck3d by’, ‘h4cked by’, ‘p4wn3d by’, ‘pwn3d by’, ‘pwnd by’, ‘pwned by’, ‘pwndz by’, ‘owned by’, ‘own3d by’, ‘touched by’, and ‘kissed by’, we consider them to be valid e.g., a message ‘This website was hacked, contact me t.me/coolhacker’ posted by a notifier ‘CoolHacker’ is considered valid. This method is looser than an exact comparison with ‘Hacked by CoolHacker’, but is still highly accurate; 100 randomly checked samples were all correct. Third, the remaining submissions are manually validated by looking for defacer signatures; some are obvious but some are complicated. Candidates are grouped by normalised handles and messages (redundant spaces removed), and then for each, the 10 most similar validated defacements are suggested to the annotator. Levenshtein distance is used to estimate the similarity between two messages, which is helpful as messages are often slightly modified from templates. If no message is found (instead images, iframes, or javascript), or leftover signatures cannot be spotted, a browser opens the defaced page to assist annotators.
@@
This assistance effectively reduces the annotation effort. One challenge is defaced pages redirecting to another page that can be modified dynamically: when they are down, the submission may point to a non-existent site, but a careful check could reveal evidence of defacers. We also consider a site is touched if its content is unchanged but the page title is modified to indicate hacking activities. We ignore cases that lack evidence to ensure those flagged valid are indeed valid. We do not use complex machine learning techniques as message texts contain lots of noise; given a small number of samples (around 10k), machine learning is not more effective than a rule-based approach. Sometimes, defacements appeared to be already verified at the collection time, but became invalid afterwards; we re-validated them months after the initial collection to make sure their status had been finalised by the archives.
@@ Appendix G Unifying Defacements and Defacers
We hashed then unified defacements across all archives based on reporting dates, original defacer handles, root victim domains, and messages left on the defaced page. Hashing reporting dates may cause repeat counts if defacers resubmit to other archives after a few days, but excluding them may lead to missing defacements of the same URL on different days due to repeat victimisation. We also unified defacers across all archives, as users tend to pick similar pseudonyms on different platforms (Goga et al., 2013). As the unification needs to be accurate and the number of unique defacers does not exceed a few thousand, machine learning is not appropriate. We instead used a semi-automatic approach combining automated handle similarity analysis with manual review. As typos may occur (e.g., missing characters, character orders, case sensitivity), similar handle pairs are first extracted using Levenshtein distance, which is set to not exceed 25% of the length. Ten messages left by defacers in each pair are then sampled to assist the annotation. A pair of handles are unified under a single nickname if their messages are closed enough, assessed based on handle appearances, message semantics, stylometry, synonyms, typos, team, nationality, languages, and handle rarity (rare ones such as ‘cj2ks’ are more likely to be used by a single person, while common ones such as ‘glory’ are more likely to be shared by multiple individuals (Liu et al., 2013)). Messages left are diverse: many defacers leave identical messages on different archives, while some are relatively similar, some are distinct and contain the defacer names, and some consist of their phone numbers. We only confirm when having sufficient evidence, uncertain pairs are left unmatched.